OpenStudy (anonymous):

find the point of intersection of the line: 3x-y-7=0 and x+5y+3+0

OpenStudy (anonymous):

r u ready 4 help?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Since you said they were line I am going to rewrite your equations first one y=3x-7 second one y=-x/5+3/5 They were just some rearrangment

OpenStudy (anonymous):

3x-7=-x/5 +3/5

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

do i multiply both sides by 5?

OpenStudy (anonymous):

you can. In fact It is a good idea

OpenStudy (anonymous):

ok for some reason I cant get the answer

OpenStudy (anonymous):

Multiplied by 5 15x-35=-x+3 16x=38 x=38/16=19/8 Plug that into either one of your equation to get Y 3x-7=y 3(19/8)-7=y

OpenStudy (anonymous):

oh ok so that is what i got...i thought it was wrong because someone said x was 2 and y was -1. THANKS